OpenStudy (anonymous):

Solve the equation: (-4) + (-1) + 2 + ------- + x = 437 Pls give steps of solution too....

OpenStudy (mathteacher1729):

What do all the dashes mean?

OpenStudy (turingtest):

\[(-4)+(-1)+2+...+x=437\]? continuation of the pattern, whatever it is? I don't get it.

OpenStudy (anonymous):

the dashes mean that there are other numbers in between 2 and x

OpenStudy (anonymous):

this loook like an A P

OpenStudy (turingtest):

well the pattern most obvious is what jimmyrep says: +3

OpenStudy (anonymous):

sum = 437

OpenStudy (anonymous):

It is an A.P. but somehow today the bell is not ringing for me!!!

OpenStudy (anonymous):

Here u r not given the terms themselves......

OpenStudy (anonymous):

437 = n/2 [-8 + (n-1)3]

OpenStudy (mathteacher1729):

This looks like the sum of an arithmetic series: http://regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

50

OpenStudy (anonymous):

that is, x = 50

OpenStudy (turingtest):

Yeah, but ow do you "show" it? I used my calculator.

OpenStudy (anonymous):

you have \[\sum_{k=0}^n3k-4=437\] there is a snappy way to sum an arithmetic series, but i always forget it. in any case you are solving for n

OpenStudy (anonymous):

yea solve for n then x = nth term

OpenStudy (anonymous):

being a bonehead and forgetting the simple way, i usually work with \[\sum_{k=1}^nk =\frac{n(n+1)}{2}\] and work from there

OpenStudy (anonymous):

lol - you're no bonehead!

OpenStudy (anonymous):

i think jimmyrep has it, the summation formula i mean

OpenStudy (anonymous):

i can work out the steps i took if you like. since i only remember that \[\sum_{k=1}^n k =\frac{n(n+1)}{2}\] i wrote the sum as \[\sum_{k=1}^n -7+3k=-7\sum_{k=1}^n1+3\sum_{k=1}^n k\] \[=-7n+\frac{3n(n-1)}{2}=437\] solve for n get \[n=19\] replace k by 19 in \[-7+3k\] get x =50

OpenStudy (anonymous):

anyway guys i'm signing - i'm just on my fourth beer and thinking is getting too hard HAPPY NEW YEAR guys

OpenStudy (anonymous):

happy new year all, see you in 2012

OpenStudy (anonymous):

Hey guys, HAPPY NEW YEAR to u ALL !!!!

OpenStudy (anonymous):

Pls check if my working below is correct...

myininaya (myininaya):

looks like i missed a cool problem

OpenStudy (turingtest):

No, you get to check Harkirat's work still :)

OpenStudy (anonymous):

The formula for the sum is Sn = n/2[2a+(n-1)d] here Sn= 437 a = -4 d = 3 so 437 = n/2[2(-4) + (n-1)(3)] 874 = n[-8 + 3n -3] which gives 3n^2 - 11n - 874 = 0 solving which we get n = 19 or n = -46/3 since term no cannot be negative, so n = 19

OpenStudy (anonymous):

so x is the 19th term using An = a + (n-1)d x = (-4) + (19-1)(3) x = -4 + 54 x = 50

OpenStudy (turingtest):

look right to me, but I had to look up that formula.

myininaya (myininaya):

(-4+0(3))+(-4+1(3))+(-4+2(3))+....+(-4+n(3))=437 (so this means x=-4+3n) \[-4(n+1)+3\sum_{i=0}^{n}i=437\] \[-4n-4+3(0)+3 \sum_{i=1}^{n}i=437\] \[-4n+3 \cdot \frac{n(n+1)}{2}=441\] \[-8n+3n(n+1)=882\] \[-8n+3n^2+3n=882\] \[3n^2-5n-882=0\] \[n=\frac{5 \pm \sqrt{25-4(3)(-882)}}{2(3)}=\frac{5 \pm 103}{6}=\frac{108}{6} \text{ or } \frac{-98}{6}\] We don't want the negative value (also that would give us an irrational answer and we want n to be an integer) so we have \[n=\frac{108}{6}=18\] => x=-4+3n=-4+3(18)=50 This is the way I would have done it. This is a nice problem! :)

OpenStudy (turingtest):

I forgot about making it -4(n+1) outside the sum sign when you start from zero, so I kept messing it up starting doing -4n. I'm a series rookie.

myininaya (myininaya):

Well I like my way because I understand it best! :) I like how Satellite and Harkirat did it different ways.

OpenStudy (anonymous):

Thanks to all!!! Actually I had to get the solution with the formulas I used. My students have not been taught the summation method/formula yet...... @TuringTest... Don,t worry, we all make mistakes. I was also making a small mistake in my calculation due to which I was getting both values of n as negative and hence was foxed. However, when I did it here, I did it right and got the positive value of n....and then it was easy to get the final answer☺

myininaya (myininaya):

Where did you get this problem?

myininaya (myininaya):

I have a problem to give to my students now. :)

myininaya (myininaya):

Can't wait!

OpenStudy (anonymous):

I take private tuition and one of the students just gave it to me over phone as he has to submit the assignment tomorrow......

myininaya (myininaya):

You do tutoring over the phone?

OpenStudy (anonymous):

*I give private tuition....

myininaya (myininaya):

private tutoring?

OpenStudy (anonymous):

Sometimes....when they forget to ask when they come to me for tuition... I even teach a kid in USA via the internet....

OpenStudy (anonymous):

Private tutoring means that besides going to school, the kids come to me for one-on-one lessons where I clear all their doubts and explain concepts as per their standard level...

myininaya (myininaya):

Cool! :)

OpenStudy (anonymous):

Are you also a teacher??☺

myininaya (myininaya):

Yes I teach.

OpenStudy (anonymous):

I am from India and you are from??

myininaya (myininaya):

United States. You know I remember you from awhile back. Its been awhile since I seen you. Do you remember me?

OpenStudy (anonymous):

Of course, I do!!!!

OpenStudy (anonymous):

Actually I have got real busy these last few months, even my nights are taken by my US student plus I got bored with solving the same type of problems again and again and do not come to openstudy so much nowadays....☻

myininaya (myininaya):

I think you always had some cool questions. I sort remember something from you a long time ago, but I can't remember it exactly.

OpenStudy (anonymous):

I guess they r interesting because I myself find them tough !!! LOL !!!! However, the moment I put them out here, somehow the solution comes running to my mind....like magic....

myininaya (myininaya):

lol

OpenStudy (anonymous):

Hey myininaya, I got a favour to ask....

myininaya (myininaya):

oh I'm sorry what?